# PP-Tac: Paper Picking Using Tactile Feedback in Dexterous Robotic Hands

### v1

@@ Abstract
Robots are increasingly envisioned as human companions, assisting with everyday tasks that often involve manipulating deformable objects. Although recent advances in robotic hardware and embodied AI have expanded their capabilities, current systems still struggle with handling thin, flat, and deformable objects such as paper and fabric. This limitation arises from the lack of suitable perception techniques for robust state estimation under diverse object appearances, as well as the absence of planning techniques for generating appropriate grasp motions. To bridge these gaps, this paper introduces PP-Tac, a robotic system for picking up paper-like objects. PP-Tac features a multi-fingered robotic hand with high-resolution omnidirectional tactile sensors R-Tac. This hardware configuration enables real-time slip detection and online frictional force control that mitigates such slips. Furthermore, grasp motion generation is achieved through a trajectory synthesis pipeline, which first constructs a dataset of finger’s pinching motions. Based on this dataset, a diffusion-based policy is trained to control the hand-arm robotic system. Experiments demonstrate that PP-Tac can effectively grasp paper-like objects of varying material, thickness, and stiffness, achieving an overall success rate of 87.5%. To our knowledge, this work is the first attempt to grasp paper-like deformable objects using a tactile dexterous hand. Our project webpage can be found at: https://peilin-666.github.io/projects/PP-Tac/.
Figure 1: Overview of the PP-Tac Robotic System. The system leverages tactile feedback of proposed round shape tactile sensor (R-Tac) in a dexterous robotic hand to pick up thin, deformable paper-like objects. (a) Hand motion generated by force and slip feedback, inspired by human paper-picking motions involving sliding and pinching. (b) Hardware setup includes a robotic arm, a dexterous hand, and four fingertip-mounted tactile sensors capable of simultaneously detecting force and slip events.
@@ I Introduction
Robots are increasingly popular as assistive agents in everyday life, particularly within household environments [38]. These robots are designed to perform various domestic tasks, often involving the grasp of thin, deformable objects such as paper and fabric [51]. For instance, clothes-folding tasks [27] require high dexterity and adaptability to accommodate variations in fabric size, texture, and stiffness, while document organization tasks [1] demand precise picking capabilities for diverse paper types and form factors. Beyond domestic settings, handling deformable objects is essential in industrial and logistical applications, such as fabricating fabrics [5] and packing objects using plastic bags and cardboard [12].
@@
Despite their significance, picking up paper-like objects remains challenging in robotics [51]. In particular, the main challenges are three-fold: 1) Vision systems, commonly used for manipulation, struggle to perceive contact information during interactions with deformable objects due to limited sensing modalities and occlusion, resulting in a lack of necessary feedback for motion planning [28]; 2) Their thin, stiff characteristics often result in flat shapes, hindering the synthesis of stable grasps using conventional methods due to insufficient contact points [9]. 3) The appearance of such objects exhibits high variability, as their shape undergoes continuous and unpredictable deformation during manipulation. These dynamic shape variations significantly impair the generalizability of vision-based methods.
@@
In contrast, humans excel at picking up paper-like objects by leveraging coordinated multi-fingered motion and tactile sensing. As shown in Fig. 1(a), the process typically begins with establishing contact using fingers, followed by sliding motions to deform the material and enable a stable pinched grasp. Such success stems from the coordination of multiple fingers, which generates the necessary friction to deform the object and utilizes sufficient finger Degree of Freedoms to adaptively establish stable contact points for a stable grasp. Additionally, tactile sensing complements visual feedback, allowing humans to perceive the object’s deformation and apply appropriate forces by detecting friction and slip. These tactile cues facilitate real-time adjustments, ensuring the successful execution of the picking-up action.
@@
Inspired by human strategies, this paper introduces a robotic system, PP-Tac (Paper-like object Picking using Tactile feedback), designed for dexterous robotic hands. The system comprises two key components: A dexterous robotic hand with hemispherical and high-resolution Vision-Based Tactile Sensors (VBTS) R-Tac. The fingertip-mounted tactile sensors provide real-time contact feedback during grasping operations. Featuring a circular sensing area and a high-frame-rate monochrome camera, this design enables faster response times and simpler calibration processes compared to conventional RGB-based tactile systems. An illustration of the system is shown in Fig. 1(b). In addition to the tactile sensor, this paper also presents A diffusion-based motion generation policy (PP-Tac policy) that imitates human picking-up skills. The proposed method first employs efficient trajectory optimization to generate expert data replicating human sliding and pinching motions. To generalize this approach to diverse deformable objects and uneven surfaces, a diffusion policy is subsequently trained using these trajectories, leveraging proprioceptive data and tactile feedback for adaptive control of the dexterous robotic hand.
@@
In comprehensive real-world experiments, the proposed PP-Tac achieved an overall success rate of 87.5% in grasping everyday thin and deformable paper-like objects, such as plastic bags, paper bags, and silk towels on flat surfaces. Fig. 1(a) illustrates examples of our arm-hand system successfully picking up paper-like objects. The PP-Tac also demonstrates significant adaptability in picking up paper-like objects on various uneven surfaces. Additionally, an ablation study further validates the contributions of each system component, highlighting the critical role of VBTS feedback and motion generation policies in achieving robust performance.
@@
    We present R-Tac, a novel circular tactile sensor designed for ease of fabrication, calibration, and scalable deployment. To demonstrate its utility, we integrate R-Tac into each fingertip of a fully actuated dexterous robotic hand, enabling real-time contact feedback during manipulation tasks
@@
    We propose a novel trajectory-optimization-based data generation framework that eliminates the need for tactile or physical simulation while achieving robust sim-to-real transfer performance.
@@
    We present the PP-Tac policy, a diffusion-based control strategy that utilizes only tactile and proprioceptive feedback for manipulating paper-like objects. This approach demonstrates robust generalization across diverse materials and surface properties.
@@
    We provide the implementation and systematic experiments of the proposed algorithms on the physical device. Both hardware and code for PP-Tac system will be open-sourced to support further research and community development.
@@ II-A Deformable Objects Manipulation
Deformable Object Manipulation (DOM) involves handling soft objects that alter shape during interaction—a ubiquitous yet challenging task in robotics. Challenges arise from uncertainties in perception and complex soft-body dynamics [16, 3, 31]. Early approaches relied on visual perception for state estimation [51, 37], enabling tasks like rope-handling [33, 37], cloth-folding [42, 27] and picking up paper with marker [14]. However, vision-based methods often underperform in real-world DOM due to varying object appearance, limited physical property perception, occlusions [25, 6], and variable lighting conditions [48, 22]. These challenges hinder the development of scalable vision-based DOM solutions for diverse environments.
Tactile sensing, particularly Vision-Based Tactile Sensors (VBTS), has demonstrated significant potential for DOM tasks compared to visual perception [51]. VBTS excel in tasks such as object shape reconstruction [34, 10, 30, 47], localization [20, 26, 7], and slip detection [44, 13], leveraging their high-resolution tactile feedback. Prior work has explored VBTS for deformable object manipulation [40], but existing implementations rely on gripper-mounted sensors, which lack the dexterity of multi-fingered hands due to limited DoF. Our experiments reveal that gripper-based approaches struggle with thin, deformable objects or those on non-flat surfaces, highlighting the need for integrating dexterous robotic hands with VBTS for robust manipulation [21].
Figure 2: The hardware design of the R-Tac and its integration into the four-fingered dexterous robotic hand system. (a) illustrates the pipeline of depth reconstruction. (b) illustrates the exploded view of the sensor, detailing each component. (c) shows the dimensions of the sensor. (d) shows the schematic design. (e) illustrates the robotic hand equipped with four sensors on its distal joint.
@@ II-B Dexterous Robotic Hand with Tactile Sensing
Current dexterous hands are often equipped with tactile sensing capabilities to enhance dexterity. Commonly used tactile sensors typically incorporate mechanisms such as capacitive [32], piezoresistive [19], or magnetic-based [15] technologies. These designs can be fabricated in various shapes and sizes, allowing them to conform to the form factor of different robotic fingers. However, the sensing principles behind these technologies limit their spatial resolution and robustness under varying environmental conditions. Although efforts have been made to develop VBTS for curved shapes [10, 46, 23, 2, 4], these three-channel imaging designs are not yet commercially available and remain challenging to deploy at scale in hand configurations. First, the illumination from tricolor light sources results in uneven intensity distributions in curved surface imaging, necessitating extensive data collection for calibration. Second, these sensors [10, 46, 23, 2] often require specialized test beds (e.g., CNC machines) to collect large datasets, increasing the calibration complexity. Third, the large data volume of three-channel images imposes bandwidth constraints, which can limit the overall frame rate in large-scale deployments. To address this, we propose R-Tac that is structurally simple, compact, easy to fabricate, and straightforward to calibrate.
Current robotic hands equipped with VBTS have been used in grasping and in-hand orientation tasks. For instance, Do et al. uses DenseTact [10], attached to an Allegro Hand, to grasp and manipulate small screws [11]. Qi et al. integrates fingertip VBTS [35] and DIGIT [43] on an Allegro Hand to enable in-hand object rotations. To the best of our knowledge, existing research has not yet explored VBTS-equipped dexterous hands for manipulating thin, deformable objects such as paper sheets.
@@ III-A Fingertip-shaped Tactile Sensing
  • High resolution: High resolution enables accurate depth reconstruction and slip detection during the picking-up process.
@@
Based on these 5 principles, the sensor design and its integration into the dexterous robotic hand is illustrated in Fig. 2. Next, we detail each component and the calibration process.
@@ III-A1 Contact and Illumination Module
The core of the sensor is a contact module (elastomer) with a uniformly illuminated, deformable sensitive surface that maintains structural rigidity during contact. Inspired by the monochrome sensing principle [29], where darkness changes indicate deformation, we developed a hemispherical structure comprising a white LED ring, a stiff transparent internal skeleton, a soft semitransparent perception layer, and a thin opaque protective layer that achieves the desired optical characteristics.
@@ III-A2 Camera Module
A micro black-and-white CMOS camera (OV9281) with a wide $160^{\circ}$ lens is used to capture the darkness data. The camera operates up to 120Hz with a resolution of $640\times 480$ and a latency of approximately 100ms.
@@ III-A4 Contact Force Estimation & Slip Detection
Our sensors are capable of detecting both contact forces and slip events. The contact force, modeled by elasticity theory, is proportional to the deformation depth and can be expressed as a function of deformation depth. The Slip Detection module is as follows:
  • Detection Model: As illustrated in Fig. 3, when the slip occurs, distinct wrinkles become visible in the sensor’s imaging. We apply a lightweight neural network architecture consisting of CNN(convolutional neural network) and MLP(multilayer perceptron) to detect the slip. The network processes a temporal sequence of the preceding five frames with a non-contact frame as input. CNN extracts the feature per image and then the feature maps are concatenated together to estimate the slip probability $P_{\text{slip}}$ through a multilayer perceptron (MLP).
  • Training: To train the network, we collected approximately 20 minutes of tactile data from the four sensors. The dataset comprises 40% slip samples and 60% non-slip samples, with each frame manually annotated to ensure ground-truth reliability. We choose binary cross-entropy as our loss function.
  • Inference: In inference, we set a threshold for $P_{\text{slip}}$. And we systematically evaluated the model’s performance by varying the decision threshold for $P_{\text{slip}}$. Empirical results demonstrate that a threshold of 0.75 yields an optimal trade-off between sensitivity and accuracy, achieving a slip detection accuracy of 86%.
Figure 3: Slip Detection. The left tactile image shows no contact, while the middle tactile image highlights wrinkle features during slip. The network computes the probability of slipping $P_{slip}$ using the no-contact tactile image and five sequential tactile images.
@@ III-B Robotic Hand System
We integrated the proposed R-Tac sensors into a fully actuated dexterous robotic hand. These tactile sensors are mounted at the distal end of each fingertip, facilitating contact characterization in the following paper-picking tasks. We designed and fabricated the robotic hand featuring 16 controllable DoFs, including the DIP, PIP, and MCP joints for the index, middle, and ring fingers, as well as the CMC, CMC-2, MCP, and IP joints for the thumb. The robotic hand is driven by Dynamixel XC330-M288-T motors, which are all multiplexed through a U2D2 Hub. For each tactile sensor, it communicates with the PC via a USB interface. The entire assembly is mounted on a Franka Research 3, a 7-DoF robotic arm, which communicates with the PC via a high-speed Ethernet connection.
@@ IV Paper-like Object Picking Problem Statement
Our approach adopts a learning-based policy rather than a model-based method due to the following challenges associated with model-based approaches:
Challenging A: series-parallel coupling. During grasping on uneven terrain, the fingers must adjust their positions vertically to conform to the surface. However, the four fingers and wrist form a coupled serial-parallel mechanism, where the motion of one finger influences the wrist state, which in turn affects the other fingers. Consequently, even if a model-based approach could predict finger positions at the next timestep, the strong dynamic coupling makes it computationally challenging to solve for all joint angles in real-time with low latency.
Challenging B: limited DoF of finger. One potential solution to the coupled serial-parallel challenge is to fix the wrist state. However, in practice, we observe that the limited DoF of the fingers prevents successful grasping with a fully constrained wrist. Intuitively, this resembles human grasping behavior: when picking up a sheet of paper from a flat surface, the wrist must first elevate and then lower to achieve stable contact.

### v2

@@ Abstract
Robots are increasingly envisioned as human companions, assisting with everyday tasks that often involve manipulating deformable objects. Although recent advances in robotic hardware and embodied AI have expanded their capabilities, current systems still struggle with handling thin, flat, and deformable objects such as paper and fabric. This limitation arises from the lack of suitable perception techniques for robust state estimation under diverse object appearances, and the absence of planning techniques for generating appropriate grasp motions. To bridge these gaps, this paper introduces PP-Tac, a robotic system for picking up paper-like objects. PP-Tac features a multi-fingered robotic hand with high-resolution round-shaped tactile sensors R-Tac for omnidirectional tactile sensing. This hardware configuration enables real-time slip detection and online force control that mitigates slips. Furthermore, grasp motion generation is achieved through a trajectory synthesis pipeline, which first constructs a dataset of fingers’ pinching motions. Then, a diffusion-based policy is trained to control the hand-arm robotic system. Experiments demonstrate that PP-Tac can effectively grasp paper-like objects of varying material, thickness, and stiffness, achieving an overall success rate of 87.5%. To the best of our knowledge, this work is the first attempt to grasp paper-like objects using a tactile dexterous hand.
Figure 1: Overview of PP-Tac. The system leverages tactile feedback from the proposed round-shaped sensor (R-Tac), integrated into a dexterous robotic hand, to grasp thin, deformable, paper-like objects. (a) Hand motions are generated by a diffusion-based policy inspired by human strategies, such as sliding and pinching. (b) The hardware setup includes a robotic arm, a dexterous hand, and four fingertip-mounted tactile sensors that simultaneously detect force and slip events.
@@ I Introduction
Robots are increasingly popular as assistive agents in everyday life, particularly within household environments [36]. These robots are designed to perform various domestic tasks, often involving the grasp of thin, deformable objects such as paper and fabric [48]. For instance, clothes-folding tasks [25] require high dexterity and adaptability to accommodate variations in fabric size, texture, and stiffness, while document organization tasks [1] demand picking capabilities for diverse paper types and form factors. Beyond domestic settings, handling deformable objects is essential in industrial and logistical applications, such as fabricating fabrics [5] and packing objects using plastic bags and cardboard [12].
@@
Despite their significance, picking up paper-like objects remains challenging in robotics [48]. In particular, the main challenges are three-fold: 1) Vision systems, commonly used for manipulation, struggle to perceive contact information during interactions with deformable objects due to limited sensing modalities and occlusion, resulting in an inaccurate environment model for motion planning [26]; 2) These objects are often flat in shape, lacking salient features for contact points and thus hindering the synthesis of stable grasps  [9]. 3) The appearance of such objects exhibits high variability, as their shape undergoes continuous and unpredictable deformation during manipulation. These dynamic shape variations significantly impair the generalizability of vision-based methods.
@@
In contrast, humans excel at picking up paper-like objects by leveraging coordinated multi-fingered motion and tactile sensing. As shown in Fig. 1 (a), the process typically begins with establishing contact with the fingers, followed by sliding motions to deform the material and to generate a contact point for the pinched grasp. Such motion is made possible by hand’s high Degree of Freedoms, which enables establishing multiple contact points adaptively during sliding motion. During this process, tactile sensing is also crucial as it allows humans to perceive the object’s deformation and decide the appropriate forces. These real-time adjustments ensure the successful execution of the picking-up action.
@@
Inspired by human strategies, this paper introduces a robotics system coined PP-Tac: Paper-like object Picking using Tactile feedback. The PP-Tac system comprises two key hardwares: A dexterous robotic hand, and the associated hemispherical and high-resolution Vision-Based Tactile Sensors (VBTS) R-Tac. The fingertip-mounted tactile sensors provide real-time contact feedback during grasping operations. Featuring a spherical sensing area and a high-frame-rate monochrome camera, this design enables faster response times and simpler calibration processes compared to conventional RGB-based tactile sensors. An illustration of the system is shown in Fig. 1(b). In addition to the tactile sensor, this paper also presents A diffusion-based motion generation policy (PP-Tac policy) that imitates human picking-up skills. The proposed method first employs efficient trajectory optimization to generate expert data replicating human sliding and pinching motions. Second, generalizability to diverse flat objects were acheived by training a diffusion policy using these trajectories, leveraging proprioceptive data and tactile feedback for adaptive control of the dexterous robotic hand.
@@
Comprehensive real-world experiments were conducted to evaluate the PP-Tac system. PP-Tac achieved an overall success rate of 87.5% in grasping everyday thin and deformable paper-like objects, including plastic bags, paper bags, and silk towels on flat surfaces. Fig. 1(a) illustrates examples of our arm-hand system successfully picking up paper-like objects. The PP-Tac also demonstrates significant adaptability in picking up paper-like objects on various uneven surfaces. Additionally, an ablation study further validated the contributions of each system component, highlighting the critical role of VBTS feedback and motion generation policies in achieving robust performance.
@@
    We present R-Tac, a novel spherical tactile sensor designed with ease of fabrication, calibration, and scalable deployment. To demonstrate its utility, we integrate R-Tac into each fingertip of a fully actuated dexterous robotic hand, enabling real-time contact feedback during manipulation tasks.
@@
    We propose a novel trajectory-optimization-based data generation framework. The proposed framework does not rely on tactile or physical simulation, which is computationally expensive, and is capable of achieving robust sim-to-real transfer.
@@
    We present the PP-Tac policy, a diffusion-based control strategy that utilizes only tactile and robot proprioceptive feedback for manipulating paper-like objects. This approach demonstrates robust generalization across diverse materials and surface properties.
@@
    We provide the implementation and systematic experiments of the proposed algorithms on a real robot system. Both hardware and code for PP-Tac system are released to support further research and community development.
@@ II-A Deformable Objects Manipulation
Deformable Object Manipulation (DOM) aims to handle soft objects that alter shape during interaction, which has been a long-standing challenging task in robotic research. Challenges mainly arised from uncertainties in perception and complex soft-body dynamics [15, 3, 29]. Early approaches solved such problem using visual perception for state estimation [48, 35], enabling tasks like rope-handling [31, 35], cloth-folding [40, 25] and picking up paper with marker [6094742]. However, vision-based methods often fall short in solving real-world DOM problems due to varying object appearance, object’s physical property that are usually unknown in advance, visual occlusions [23, 6], and variable lighting conditions [46, 20]. These challenges hinder the development of scalable vision-based DOM solutions for diverse environments.
Tactile sensing, particularly Vision-Based Tactile Sensors (VBTS), has demonstrated significant potential for solving DOM tasks [48]. Leveraging their high-resolution tactile feedback, VBTS has demonstrated high performance in object shape reconstruction [32, 10, 28, 45], localization [18, 24, 7], and slip detection [42, 13]. Prior work has explored VBTS for deformable object manipulation [38], but existing implementations rely on gripper-mounted sensors, which lack the dexterity of multi-fingered hands due to limited DoF. Our experiments reveal that gripper-based approaches struggle with thin deformable objects, and lack sufficient adapability for objects placed on non-flat surfaces, which highlights the need for integrating dexterous robotic hands with VBTS for robust manipulation [19].
Figure 2: The hardware design of the R-Tac and its integration into the four-fingered dexterous robotic hand system. (a) illustrates the pipeline of depth reconstruction. (b) illustrates the exploded view of the sensor, detailing each component. (c) shows the dimensions of the sensor. (d) shows the schematic design. (e) illustrates the robotic hand equipped with four sensors as its distal links.
@@ II-B Dexterous Robotic Hand with Tactile Sensing
Current dexterous hands are often equipped with tactile sensors. Commonly used tactile sensors typically incorporate mechanisms such as capacitive [30], piezoresistive [17], or magnetic-based [14] technologies. These designs can be fabricated in various shapes and sizes, allowing them to conform to the form factor of different robotic fingers. However, the sensing principles behind these technologies limit their spatial resolution and robustness under varying environmental conditions. To enhance sensing quality, recent work has devoted efforts to develop various VBTS sensors, especially with curved elastomer surfaces [10, 44, 21, 2, 4]. However, most of these VBTS sensors are not yet commercially available, and remain challenging to deploy at scale in hand. This is mainly due to the challenge arised from sensor’s calibration, as the illumination from RGB chromatic light sources results in uneven light intensity distributions on curved elastomer surface, necessitating extensive data collection for calibration. During such data collection, these sensors [10, 44, 21, 2] often require specialized test beds (e.g., fabricated using CNC machines) to collect large datasets, increasing the calibration complexity. Moreover, transmitting realtime chromatic video streams imposes higher bandwidth requirements, which can limit the overall frame rate in large-scale deployments. To address the above issues, we propose R-Tac that is structurally simple to fabricate, compact, and easy to calibrate.
Current robotic hands equipped with VBTS have been used in grasping and in-hand orientation tasks. For instance, Do et al. uses DenseTact [10] attached to an Allegro Hand, to grasp and manipulate small screws [11]. Qi et al. integrates fingertip VBTS [33] and DIGIT [41] on an Allegro Hand to rotate objects in hand. To the best of our knowledge, existing research has not yet explored VBTS-equipped dexterous hands for manipulating thin, deformable objects such as paper sheets.
@@ III-A Fingertip-shaped Tactile Sensing
  • High resolution: High resolution enables accurate depth reconstruction and slip detection during picking-up.
@@
Based on these 5 principles, the sensor design and its integration into the dexterous hand is illustrated in Fig. 2. Next, we detail each component and the calibration process.
@@ III-A1 Contact and Illumination Module
The core of the sensor is a contact module (elastomer) with a uniformly illuminated, deformable sensitive surface that maintains structural rigidity during contact. Inspired by the monochrome sensing principle [27], where intensity changes indicate deformation, we developed a hemispherical structure comprising a white LED ring, a stiff transparent internal skeleton, a soft semitransparent perception layer, and a thin opaque protective layer that achieves the desired optical characteristics.
@@ III-A2 Camera Module
A micro black-and-white CMOS camera (OV9281) with a wide $160^{\circ}$ lens is used to capture the light intensity data. The camera operates up to 120Hz with a resolution of $640\times 480$ and a latency of approximately 100ms.
@@ III-A4 Contact Force Estimation & Slip Detection
Our sensors are capable of detecting both contact forces and slip events. The contact force, modeled by elasticity theory, is proportional to the deformation depth as a linear function. The slip detection module is as follows:
  • Detection Model: As illustrated in Fig. 3, when the slip occurs, distinct wrinkles become visible in the sensor’s imaging. We apply a lightweight neural network architecture consisting of CNN (convolutional neural network) and MLP (multilayer perceptron) to detect the slip. The network processes a temporal sequence of the preceding five frames with a non-contact frame as input. CNN extracts the feature per image and then the feature maps are concatenated together to estimate the slip probability $P_{\text{slip}}$ through a MLP.
  • Training: To train the network, we collected approximately 20 minutes of tactile data from four sensors. The dataset comprises 40% slip samples and 60% non-slip samples, with each frame manually annotated. We choose binary cross-entropy as our loss function.
  • Inference: In inference, it is necessary to set a threshold for $P_{\text{slip}}$ for positive detection. Through manual adjustments, empirical results demonstrate that a threshold of 0.75 yields an optimal trade-off between sensitivity and accuracy, achieving a slip detection accuracy of 86%.
Figure 3: Slip detection. The left tactile image shows no contact, while the middle tactile image highlights wrinkle features during slip. The network computes the probability of slipping $P_{slip}$ using the no-contact tactile image and five most recent tactile images.
@@ III-B Robotic Hand System
We integrated the proposed R-Tac sensors into a fully actuated dexterous robotic hand. These tactile sensors are mounted at the distal end of each fingertip, facilitating contact characterization in the following paper-picking tasks. We designed and fabricated the robotic hand featuring 16 controllable DoFs, including the DIP, PIP, and MCP, MCP-2 joints for the index, middle, and ring fingers, as well as the CMC, CMC-2, MCP, and IP joints for the thumb. The robotic hand is driven by Dynamixel XC330-M288-T motors, which are all multiplexed through a U2D2 Hub. For each tactile sensor, it communicates with the PC via a USB interface. The entire assembly is mounted on a Franka Research 3, a 7-DoF robotic arm, which communicates with the PC via a high-speed Ethernet connection.
@@ IV Paper-like Object Picking Problem Statement
One challenge is determining the control inputs for all joints and the wrist pose. Intuitively, this resembles human grasping behavior: when picking up a sheet of paper from a flat surface, the wrist must first elevate and then lower to establish stable contact. However, a finger-wrist coupling issue arises: the motion of one finger requires a specific wrist state, which in turn affects the movement of other fingers. In practice, our approach solved this problem by adopting a learning-based policy rather than a model-based optimization paradigm. This is due to its superior efficiency in deployment, as we found model-based optimization is too computationally expensive to adapt for online execution.